Find the equation for the tangent plane to the surface x-z= 4 arctan(yz) at the point (1 +pi,1,1).
1
Expert's answer
2020-10-05T18:20:34-0400

We have "x-4arctan(yz)-z=0" and point "(1+\pi, 1, 1)."

The equation is "F_{x}^{'}(x_0, y_0, z_0)(x-x_0)+F_{y}^{'}(x_0, y_0, z_0)(y-y_0)+F_{z}^{'}(x_0, y_0, z_0)(z-z_0)=0."

"F_{x}^{'}=1", "F_{x}^{'}(1+\pi, 1, 1)=1."

"F_{y}^{'}=-4\frac{1}{y^2z^2+1}z=\frac{-4z}{y^2z^2+1}", "F_{y}^{'}(1+\pi, 1, 1)=\frac{-4}{2}=-2".

"F_{z}^{'}=-4\frac{1}{y^2z^2+1}y-1=\frac{-4y}{y^2z^2+1}-1", "F_{z}^{'}(1+\pi, 1, 1)=\frac{-4}{2}-1=-3".

So, "1(x-1-\pi)-2(y-1)-3(z-1)=0",

"x-1-\pi-2y+2-3z+3=0",

"x-2y-3z+4-\pi=0".


Answer: "x-2y-3z+4-\pi=0."